OpenStudy (anonymous):

Hi, I need help with calculus. How do I write an exponential function when they give me two coordinates on a graph? -I Appreciate the help!

zepdrix (zepdrix):

Can you provide an example? What coordinates did they give you? :) That will make it a bit easier to explain.

OpenStudy (anonymous):

They gave me (0,2) and (2,5)

zepdrix (zepdrix):

Like I said, the numbers came out very awkward looking :( I don't think I made a mistake in there... hmm..

OpenStudy (anonymous):

hmmm ok. But in the textbook it says to find the exponential function y= Y(subscribe 0) e^kt. Is that the samething?

zepdrix (zepdrix):

\(y_o=C\). But I didn't allow for the possibility of a \(k\) value. My mistake, that's why the numbers are coming out all nasty. Let's try that again! :)

zepdrix (zepdrix):

So the instructions are telling you to use the base \(e\) for this exponential function?

OpenStudy (anonymous):

yes. " Find the exponential function y=Y(subscribe 0) e^kt whose graphs passes through the two points"

zepdrix (zepdrix):

\[\large y=y_o e^{kt}\] Plugging in our first coordinate (0,2) gives us,\[\large 2=y_oe^0\qquad \rightarrow\qquad y_o=2\] Giving us,\[\large y=2e^{kt}\]Plugging in our other coordinate (2,5),\[\large 5=2e^{2k}\]To solve for k, we'll have to do a little algebra? Remember your rules of logarithms? :) That's what we'll need to apply right now.

zepdrix (zepdrix):

Let's start by dividing both sides by 2, then taking the natural log of both sides.\[\large \ln\left(\frac{5}{2}\right)=\ln\left(e^{2k}\right)\] Recalling an important rule of logarithms,\[\large \log\left(a^b\right)=b \cdot \log(a)\]We can bring the exponent out as a coefficient in front of the log. Applying that gives us,\[\large \ln\left(\frac{5}{2}\right)=2k \cdot \ln\left(e\right)\] The natural log of e is just 1. \[\large 2k=\ln\left(\frac{5}{2}\right)\] With just a little bit more work, you should be able to find your \(k\) value! :)

zepdrix (zepdrix):

Do you by chance have an answer key that we can check this against? :D

OpenStudy (anonymous):

Yes. That's what the textbook has an answer. Thanks.